Accelerator Update for June 2, 2008 - June 4, 2008
Monday June 2
The day shift began with Operations monitoring Tevatron (TeV) store 6185 (63.28E30) and the Recycler stash (325.2E10), with the Antiproton Source (Pbar) stacking (33.3E10), with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the NuMI and SciBooNE / MiniBooNE (BNB) experiments taking beam.
FESS personnel reported at 8:15 AM that they were beginning work on the MINOS sump #3. They completed the job later in the afternoon.
Operations terminated store 6185 at 10:16 AM.
An RF expert began work on a TeV RF station (TRF8) at 10:46 AM. Later, the expert reported that he would be replacing the driver tubes on TRF4 & 8.
At 11:07 AM, a Booster vacuum pump (IPRF23) tripped off and took down BRF13.
At 11:08 AM, the Booster West Anode Power Supply (WAPS) tripped off on a modulator door open and brought down all the Booster RF stations. Everything, including BRF13 was back up and running by 11:18 AM.
At 12:04 PM, KRF3 tripped off. It reset.
At 12:36 PM, KRF4 tripped off. It reset.
At 12:50 PM, LRF5 tripped off. It reset.
Operations reported at 12:55 PM that the TeV RF system was ready for shot setup.
At 1:23 PM, TRF1 tripped off. An expert investigated and had it back on 18 minutes later.
Operations reported at 1:41 PM that the C4 cryo system was ready to go.
At 1:56 PM, KRF3 tripped off. It reset.
Operations reported at 2:09 PM that water technicians were flushing the TeV sector B0 heat exchanger.
Operations put the TeV through a dry squeeze at 2:10 PM.
Operations put the TeV into shot setup at 2:39 PM.
NuMI experts filled the horn #1 LCW system. Operations resumed sending beam to NuMI at 3:59 PM.
Operations reported at 4 PM that the Recycler had stashed antiprotons once during the day shift.
Operations established store 6187 at 4:26 PM with an initial luminosity of 254.26E30.
At 5 PM, Operations turned off the NuMI beam so experts could add more water to the horn cooling system.
Operations resumed sending beam to NuMI at 5:15 PM.
At 5:33 PM, LRF2, 3, and 5 all tripped off. Everything reset.
At 5:38 PM, Operations turned off the NuMI beam; the LCW fill did not take. NuMI experts reported that horn #1 would have to be replaced. The preliminary work will begin on 6/3/08.
At 5:39 PM, LRF5 tripped off. It reset.
At 5:43 PM, KRF7 tripped off. It reset.
At 5:47 PM, LRF4 tripped off. It reset.
At 5:49 PM, LRF1, 2, 3, and 5 all tripped off. A Linac expert investigated. Linac resumed sending beam at 6:38 PM.
Recycler experts reported at 6:44 PM that they were turning off the VME dampers.
At 7:57 PM, a Booster kicker (MKS06) tripped off and would not reset. Operators adjusted the four remaining kickers to compensate. Operations contacted experts.
Cryo system experts at CHL reported at 11:57 PM that their screw compressor B failed. They started up the backup compressor.
Operations reported at midnight that the Recycler had stashed four times during the evening shift.
Tuesday June 3
The midnight shift began with Operations monitoring store 6187 (101.65E30) and the stash (217.3E10), with Pbar stacking (40.0E10), and with BNB taking beam. NuMI was off for horn replacement, which will take three to four weeks.
At 12:14 AM, KRF7 tripped off. It reset.
At 12:42 AM, the MI RF system tripped off due to its South Anode Power Supply tripping off. Operations contacted experts who helped operators turn everything back on.
Pbar resumed stacking at 12:50 AM.
At 12:53 AM, BRF11 tripped off. It reset.
At 1:48 AM, Operations reported that the Booster GMPS #2 had tripped off four times in a row on a filter current indication. The Booster RF was sparking excessively. Operations contacted an expert who told operators to bypass GMPS #2 and bring GMPS #4 back online.
Booster resumed sending beam to Pbar and BNB at 2:43 AM.
At 3:12 AM, Operations reported that while Pbar was in standby, during the Booster work, Pbar began loosing some antiprotons from the core and lifetime went down. An operator tuned on stacktail cooling and pushed the beam back into the core. Operations informed an expert.
A Pbar expert reported at 6:19 AM that Pbar had set a New Stacking Record with a rate of 27.01 mA per hour.
Operations terminated store 6187 at 6:57 AM.
Operations put the TeV into shot setup at 7:09 AM.
Operations reported at 7:53 AM that they were ready to load protons into the TeV, but CHL reported that it would be another 45 minutes before compressor C was online.
Operations reported at 7:54 AM that the Recycler had stashed antiprotons five times during the midnight shift.
At 8:09 AM, a MI kicker (MPS5S) was alarming on a “Current Limit” indication, but the kicker was still extracting beam. Operators investigated.
Operations turned off Pbar at 8:23 AM in preparation for an access.
Operations turned off the Booster beam at 8:32 AM so a kicker expert could investigate MKS7.
The picture on shows the very beginning of Fermilab’s particle accelerators: the Cockcroft Walton.
The high voltage dome is kept at a -750 thousand volts (kV).
The tube is made of spun glass.
The tube contains the Ion Source and the Accelerating Column. The column is a pressure vessel that also contains an insulating gas to reduce the chances of sparking. The column consists of seven discs that guide the ion beam.
Two water resistors pass from the dome through the outer edges of rings, which surround the column. Their purpose is to maintain an equal voltage potential difference between the discs.
A Linac expert reported at 8:37 AM that he was looking at a quadrupole power supply (QPS103).
A Preaccelerator expert reported at 8:44 AM that the I- Source generator brushes were bad and needed replacing.
Linac experts began work on KRF4 at 8:47 AM.
The Pbar AP0 water cage access began at 8:48 AM.
Recycler experts began replacing the damper at 8:57 AM.
Operations established store 6188 at 9:11 AM with an initial luminosity of 246.03E30.
Preaccelerator experts reported at 9:24 AM that they had completed the I- Source water resistor flush and the generator brush change.
A RF expert began work on MIRF11 at 9:25 AM. He was done about 14 minutes later.
Operations reported that all Linac work had been completed. Operators established beam to the Linac dump at 10:18 AM.
Operations established beam to BNB at 10:29 AM.
Pbar resumed stacking at 10:34 AM.
At 11:13 AM, the 400 MeV area beam permit tripped off. It reset.
At 1:59 PM, LRF1 tripped off. It reset.
At 2:54 PM, MIRF3 tripped off. An expert investigated. He had the station up five minutes later.
Operations reported at 4 PM that the Recycler had stashed five times during the day shift.
At 7:35 PM, KRF6 tripped off. It reset.
At 9:14 PM, KRF6 tripped off and would not reset. After numerous attempts, operators contacted an expert who suggested they adjust the switch balance. This worked. Linac resumed sending beam at 9:58 PM.
At 11:12 PM, KRF6 tripped off again. Operators adjusted the switch balance and reset the station.
Operations reported at 11:59 PM that the Recycler had stashed five times during the evening shift.
Wednesday June 4
The midnight shift began with Operations monitoring store 6188 (64.45E30) and the stash (316.3E10), with Pbar stacking (29.6E10), and with BNB taking beam.
Operations terminated store 6188 at 1:38 AM.
At 1:45 AM, KRF6 tripped off. It reset.
At 1:48 AM, cryo system experts at CHL reported that the TeV sector C3 wet engine might be failing. They will investigate.
At 2:26 AM, cryo experts reported that a piston on the TeV sector C3 wet engine needed adjusting. They put the cryo system into JT valve mode and called in technicians.
Operators reported at 2:33 AM that they had replaced the TeV C:HD12 corrector magnet power supply.
At 3:02 AM, LRF2 tripped off. It reset.
Operations put the TeV through a dry squeeze at 3:14 AM and at 5:16 AM.
Operations put the TeV into shot setup at 5:43 AM.
At 6:30 AM, two MI kickers (I:KPS4A & B) tripped off. They both reset.
Operators reported that they had the Pelletron ready for access at 7:42 AM.
Operations established store 6191 at 7:51 AM with an initial luminosity of 258.28E30.
Operations reported at 8:13 AM that the Recycler had stashed two times during the midnight shift.
The Plans for Wednesday
The plans for today are to stack, stash, and store.
Comments and Suggestions
|last modified 06/09/2008 email Fermilab|
|Security, Privacy, Legal|